x/3 = x/2 -2
I cant remember how to do this.
Thanks,
Pat

Answer by stanbon(75887)

x/3 = x/2 -2
Multiply through by 6 to get:
2x=3x-12
-x=12
x=-12
Cheers,
Stan H.